a cracker of mass 100 g explodes into 2 equal mass . show that these 2 pieces of cracker flies inopposite direction
1
Expert's answer
2012-09-28T09:05:45-0400
In a closed system (one that does not exchange any matter with the outside
and is not acted on by outside forces) the total momentum P =const.
P = M
* V0, where M = 100 g, V0 = 0. So, P = 0.
After the explosion, P = m1 * V1 +
m2 * V2, where m1 = m2 = M / 2.
(M / 2) * V1 + (M / 2) * V2 = 0
V1 + V2 =
0
V1 = - V2, so these 2 pieces of cracker should fly in opposite directions.